OpenStudy (anonymous):

HELP WILL FAN AND GOLD Write an equation of a line that passes through 9-2,-1) with slope of 3

OpenStudy (anonymous):

(-2,-1) *

OpenStudy (anonymous):

You have a point in the form (x1,y1) and a slope m so just use \[y - y1 = m(x - x1)\] and solve for y

OpenStudy (anonymous):

can you walk me through it, i'm so unbelievably bad at slope you don't even know

OpenStudy (anonymous):

Sure. So first thing to do is to plug in all your numbers into the equation. I've labeled everything so just replace letters with numbers and write what you get on here :)

OpenStudy (anonymous):

so it would be y-2=3(x-1)?? I know that's probably wrong lol

OpenStudy (anonymous):

You have the right idea but the only thing wrong is your negatives. Remember you have a point (-2,-1) so what you really should do is \[y - (-2) = 3(x - (-1))\] now you're minusing a negative number so that turns it into a positive number therefore you would get \[y + 2 = 3(x + 1)\] you were very close, but you just gotta be careful of your negatives! :)

OpenStudy (anonymous):

I knew it was something like that! While i've got you one here do you think you can help with another one?

OpenStudy (anonymous):

Then all you have to do now is just solve for y. And sure

OpenStudy (anonymous):

Okay the question is Use the graph to determine how many solutions exist for the system -4x+3y=12 and x+y=2

OpenStudy (anonymous):

A)0 B)1 C)2 D) infinitely many

OpenStudy (anonymous):

Give me just one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does it give you a graph? Or does it want you to graph them yourself?

OpenStudy (anonymous):

yes it gives you a graph it has to lines intersecting

OpenStudy (anonymous):

How many intersections are there? The number of intersections equals the number of solutions :)

OpenStudy (anonymous):

1 intersection, so it be 1! ahh that was so easy I just forgot about the whole thing

OpenStudy (anonymous):

Yep! Easy peasy! :)